Region 1995 1997 2000
Northe 51.4 51.6 53.6
Midwest 61.8 62.5 64.4
South 91.8 94.2 100.2
West 57.7 59.4 63.2
find probability that a US resident selected random satisfy
a.lived in west in 1997
b.lived in midwest in 1995
c.lived in northe or midwest in 1997
d.what are the odds that a randomly selected U.S resident in 2000 was not from the south.

Answer by checkley77(12844)

Region 1995 1997 2000
Northe 51.4 51.6 53.6
Midwest 61.8 62.5 64.4
South 91.8 94.2 100.2
West 57.7 59.4 63.2
West in 1997:
59.4/267.7=.2219 or 22.19% probability they lived in the West.
Midwest in 1995:
61.8/262.7.2352 or 23.52%
North or Midwest in 1997:
114.1/267.7=.4262 or 42.62%.
Not from the South in 2000:
(281.4-100.2)/281.4=181.2/281.4=.6439 or 64.39%